OpenStudy (anonymous):

I will medal!!! Evaluate the summation of negative 5 n minus 1, from n equals 3 to 12. Answers: -410 -385 385 410

OpenStudy (aihberkhan):

I actually found your question on another open study post! Why don't you take a look at that and see if that helps you! http://openstudy.com/updates/53f9ff59e4b0010fcd983928 :)

OpenStudy (anonymous):

I dont understand how they got that though and he doesn't tell the answer @AihberKhan

OpenStudy (aihberkhan):

Okay. All they are saying is that you have to solve all of these: -5(3)-1 = -5(4)-1 = -5(5)-1 = -5(6)-1 = -5(7)-1 = -5(8)-1 = -5(9)-1 = -5(10)-1 = -5(11)-1 = -5(12)-1 = and then add up the answers for each of them. @Anastasia2

OpenStudy (anonymous):

so...what is it then? i got 385 @AihberKhan

OpenStudy (aihberkhan):

Well tell me what you got for each of them.